# Parrondo’s effects with aperiodic protocols

### v1

@@ Abstract
In this work, we study the effectiveness of employing archetypal aperiodic sequencing – namely Fibonacci, Thue-Morse, and Rudin-Saphiro – on the Parrondian effect. From a capital gain perspective, our results show that these series do yield a Parrondo’s Paradox with the Thue-Morse based strategy outperforming not only the other two aperiodic strategies but benchmark Parrondian games with random and periodical ($AABBAABB\ldots$) switching as well. The least performing of the three aperiodic strategies is the Rudin-Shapiro. To elucidate the underlying causes of these results, we analyze the cross-correlation between the capital generated by the switching protocols and that of the isolated losing games. This analysis reveals that a pronounced anti-correlation (below -0.95) with both isolated games is typically required to achieve a robust manifestation of Parrondo’s effect. We also study the influence of the sequencing on the capital using the lacunarity and persistence measures. In general, we observe that the switching protocols tend to become less performing in terms of the capital as one increases the persistence and thus approaches the features of an isolated losing game. For the (log-)lacunarity, a property related to heterogeneity, we notice that for small persistence (less than 0.5) the performance increases with the lacunarity with a maximum around 0.4. In respect of this, our work shows that the optimisation of a switching protocol is strongly dependent on a fine-tuning between persistence and heterogeneity.
@@ I Introduction
From basic problems such as Archimedean hydrostatics to more sophisticated cases of special relativity, Physics and other sciences are full of counter-intuitive results111 These counter-intuitive phenomena are often called paradoxes. In this case, the term paradox is ‘veridical’, ie., one whose ‘proposition’ or conclusion is in fact true despite its air of absurdity and not a ‘falsidical’ paradoxQuine (1976). . Along these lines, ratchets turned into a trendy topic in Statistical Physics and related fields Feynman, Leighton, and Sands (1963); Parrondo and Español (1996) when ‘paradoxes’ are the subject, because of its relation to the capacity to perform useful work. That especially includes the case of the so-called flashing ratchetProst et al. (1994), where a periodic potential in space (eg, sawtooth profile) is switched on and off. With that, one manages to impose a drift – ie, to perform a net useful work – to a Brownian particle even though the potential presents an overall zero gradient. It was in this framework that J.M.R. Parrondo devised a combination of losing games into an advantageous strategy that switches between the losing gamesParrondo (1996).
Despite the fact that similar counter-intuitive concepts were also conveyed in other subjectsAbbott (2001) such as biophysicsWesterhoff et al. (1986), nonlinear dynamicsKey (1987), granular matterRosato et al. (1987), random diffusionPinsky and Scheutzow (1992), biochemistryAjdari and Prost (1992), financeMaslov and Zhang (1998); Luenberger (1998), and ecologyJansen and Yoshimura (1998) before the publication of the Parrondo’s effect itself, the model has been pivotal in helping understand how the stochastic thermal fluctuations or other sources of fluctuation and disorder – eg, imperfections in electronic devicesBucolo et al. (2021) – actually help achieve a positive outcome in a variety of systems ranging from exact and natural sciences to sociology and engineeringRosato et al. (1987); Pinsky and Scheutzow (1992); Heath, Kinderlehrer, and Kowalczyk (2002); Koh and Cheong (2018); Osipovitch, Barratt, and Schwartz (2009); Cheong et al. (2016); Cheong, Koh, and Jones (2019); Wen and Cheong (2024). Because of their conceptual association with the flashing ratchet mechanism, Parrondian systems have been studied paying particular attention to the impact of the periodicity or quasi-periodicityRech, Beims, and Gallas (2007) in which the two losing games, $A$ and $B$, are combined to establish regular sequences of $A$s and $B$s or purely random schemes; midway, we have aperiodic sequencesLind, Corte-Real, and Gallas (2001) that are obtained by the application of specific deterministic rules, but for which it is not possible to establish a regular patternLynn (1989). Given the widespread usefulness of aperiodic sequences it is worth testing whether combinations of losing games ruled by that sort of sequencing can yield Parrondian systems. This is the main goal of the present work.
@@ V Final remarks
The RS generator is the case in which the capital depicts the roughest curve that makes it outperform the random switching protocol during some periods and underperform the latter in others. Actually, for $t=10^{4}$ it is not possible to distinguish the distribution of capital between RS and the random protocol with statistical significance [^2] The same occurs for the canonical quintessential periodical Parrondo’s gameHarmer and Abbott (1999). Bearing in mind that the autocorrelation function of the RS sequence is equal to that of a white noise, this result is understandable. It is worth noting that although the white noise correlation function, the application of non-linear measures (e.g. Lempel–Ziv complexity) shows that the RS sequence is not a purely randomPires and Queirós (2020); Pires, Crokidakis, and Duarte Queirós (2022).
@@
In order to shed further light on the relation between the features of the Parrondian game and its effectiveness, we have carried out the analysis looking at the cross-correlation between the capital of a switching protocol and its underlying games. For our aperiodic protocols, we have found the cross-correlations between the capital and each individual game are strongly negative, an unsurprising result since each isolated game is a losing strategy and the protocol combining them is a winning one. As a matter of fact, we expect the more anticorrelated the outcome of the switching protocol is with that of each isolated game, the more performing the former. This observation is best understood when we combine the cross-correlation Fig. 4 with the capital gain results presented in Fig. 5. All the three aperiodic switching protocol present very high anti-correlation with isolated games with TM being the most anti-correlated of them and the RS the very slightest less anti-correlated of them. [^3] For the sake of simplicity, when we mention the anti/cross-correlation between strategies and games it must be read as the anti/cross-correlation between the capital of the switching protocol and that of each isolated game. On the other hand, we verify that the P[1,1] (an anticorrelated protocol) is strongly correlated to both games, a trait that is related the poor performance of the combined game. Other periodical strategies, namely P[1,3], P[3,3], and P[4,3] go along this reasoning. This relation between cross-correlation and performance is verified for the majority of the cases, but it is neither universal nor univocal though, which makes it a future subject of study.
@@
Still with the goal of understanding the connection between the performance of a switching protocol with the properties of its sequencing scheme, we have introduced a structural plane composed of measurements of lacunarity and persistence analogous to the complexity-entropy plane used to classify data seriesRosso et al. (2007). Overall, we observe that the switching protocols are prone to become less performing as its persistence augments. This is understood by the fact that highly persistent strategies tend to look very much like an isolated game. Yet, there is the heterogeneity assessed by (log-)lacunarity, which can dramatically change the performance. For instance, the TM switching protocol outperforms both the RS and the periodical P[3,2]. These three cases have close log-lacunarity values, but significantly larger persistence one with respect to the other; however, when we compare the performance of P[4,1] with P[1,4], which have the same persistence, we understand that the increase of lacunarity is accompanied by a significant capital hike. A close picture is learned when we go from the Fb game to the periodic P[2,1] and therefrom to the TM game or farther afield to P[1,2]. As well-known, a strategy becomes Parrondian when the second game is played at the right time leading to a positive accumulation of capital; the lacunarity is a way to gauge that sparseness with respect to one of the games (ie, game $A$ in our case), an insight that the cross-correlation is unable to provide us with. This explains the reason game P[1,4] has high lacunarity (wherein game $A$ is less frequent) than game P[4,1], for which game $A$ is rather frequent. In future work, we expect to explore the delicate balance between lacunarity and persistence from a analytical point of view as well as surveying the performance of these aperiodic series in history-dependent switching protocols.

### v2

@@ Abstract
In this work, we study the effectiveness of employing archetypal aperiodic sequencing – namely Fibonacci, Thue-Morse, and Rudin-Shapiro – on the Parrondian effect. From a capital gain perspective, our results show that these series do yield a Parrondo’s Paradox with the Thue-Morse based strategy outperforming not only the other two aperiodic strategies but benchmark Parrondian games with random and periodical ($AABBAABB\ldots$) switching as well. The least performing of the three aperiodic strategies is the Rudin-Shapiro. To elucidate the underlying causes of these results, we analyze the cross-correlation between the capital generated by the switching protocols and that of the isolated losing games. This analysis reveals that a strong anticorrelation with both isolated games is typically required to achieve a robust manifestation of Parrondo’s effect. We also study the influence of the sequencing on the capital using the lacunarity and persistence measures. In general, we observe that the switching protocols tend to become less performing in terms of the capital as one increases the persistence and thus approaches the features of an isolated losing game. For the (log-)lacunarity, a property related to heterogeneity, we notice that for small persistence (less than 0.5) the performance increases with the lacunarity with a maximum around 0.4. In respect of this, our work shows that the optimisation of a switching protocol is strongly dependent on a fine-tuning between persistence and heterogeneity.
https://pubs.aip.org/aip/cha/article-abstract/34/12/123126/3323833/Parrondo-s-effects-with-aperiodic-protocols?redirectedFrom=fulltext
@@ I Introduction
From basic problems such as Archimedean hydrostatics to more sophisticated cases of special relativity, Physics and other sciences are full of counter-intuitive results111 These counter-intuitive phenomena are often called paradoxes. In this case, the term paradox is ‘veridical’, i.e., one whose ‘proposition’ or conclusion is in fact true despite its air of absurdity and not a ‘falsidical’ paradoxQuine (1976). . Along these lines, ratchets turned into a trendy topic in Statistical Physics and related fieldsFeynman, Leighton, and Sands (1963); Parrondo and Español (1996) when ‘paradoxes’ are the subject, because of its relation to the capacity to perform useful work. That especially includes the case of the so-called flashing ratchetProst et al. (1994), where a periodic potential in space (e.g., sawtooth profile) is switched on and off. With that, one manages to impose a drift – i.e., to perform a net useful work – to a Brownian particle even though the potential presents an overall zero gradient. It was in this framework that J.M.R. Parrondo devised a combination of losing games into an advantageous strategy that switches between the losing gamesParrondo (1996).
Despite the fact that similar counter-intuitive concepts were also conveyed in other subjectsAbbott (2001) such as biophysicsWesterhoff et al. (1986), nonlinear dynamicsKey (1987), granular matterRosato et al. (1987), random diffusionPinsky and Scheutzow (1992), biochemistryAjdari and Prost (1992), financeMaslov and Zhang (1998); Luenberger (1998), and ecologyJansen and Yoshimura (1998) before the publication of the Parrondo’s effect itself, the model has been pivotal in helping understand how the stochastic thermal fluctuations or other sources of fluctuation and disorder – e.g., imperfections in electronic devicesBucolo et al. (2021) – actually help achieve a positive outcome in a variety of systems ranging from exact and natural sciences to sociology and engineeringRosato et al. (1987); Pinsky and Scheutzow (1992); Heath, Kinderlehrer, and Kowalczyk (2002); Koh and Cheong (2018); Osipovitch, Barratt, and Schwartz (2009); Cheong et al. (2016); Cheong, Koh, and Jones (2019); Wen and Cheong (2024). Because of their conceptual association with the flashing ratchet mechanism, Parrondian systems have been studied paying particular attention to the impact of the periodicity or quasi-periodicityRech, Beims, and Gallas (2007) in which the two losing games, $A$ and $B$, are combined to establish regular sequences of $A$s and $B$s or purely random schemes; midway, we have aperiodic sequencesLind, Corte-Real, and Gallas (2001) that are obtained by the application of specific deterministic rules, but for which it is not possible to establish a regular patternLynn (1989). Given the widespread usefulness of aperiodic sequences it is worth testing whether combinations of losing games ruled by that sort of sequencing can yield Parrondian systems. This is the main goal of the present work.
@@ V Final remarks
The RS generator is the case in which the capital depicts the roughest curve that makes it outperform the random switching protocol during some periods and underperform the latter in others. Actually, for $t=10^{4}$ it is not possible to distinguish the distribution of capital between RS and the random protocol with statistical significance [^2] The same occurs for the canonical quintessential periodical Parrondo’s gameHarmer and Abbott (1999). Bearing in mind that the autocorrelation function of the RS sequence is equal to that of a white noise, this result is understandable. It is worth noting that although the white noise correlation function, the application of non-linear indicators (e.g. Lempel–Ziv measure) shows that the RS sequence is not a purely randomPires and Queirós (2020); Pires, Crokidakis, and Duarte Queirós (2022).
@@
In order to shed further light on the relation between the features of the Parrondian game and its effectiveness, we have carried out the analysis looking at the cross-correlation between the capital of a switching protocol and its underlying games. For our aperiodic protocols, we have found the cross-correlations between the capital and each individual game are strongly negative, an unsurprising result since each isolated game is a losing strategy and the protocol combining them is a winning one. As a matter of fact, we expect the more anticorrelated the outcome of the switching protocol is with that of each isolated game, the more performing the former. This observation is best understood when we combine the cross-correlation Fig. 4 with the capital gain results presented in Fig. 5. All the three aperiodic switching protocol present very high anticorrelation with isolated games with TM being the most anticorrelated of them and the RS the very slightest less anticorrelated of them [^3] For the sake of simplicity, when we mention the anti/cross-correlation between strategies and games it must be read as the anti/cross-correlation between the capital of the switching protocol and that of each isolated game.. On the other hand, we verify that the P[1,1] (an anticorrelated protocol) is strongly correlated to both games, a trait that is related the poor performance of the combined game. Other periodical strategies, namely P[1,3], P[3,3], and P[4,3] go along this reasoning. This relation between cross-correlation and performance is verified for the majority of the cases, but it is neither universal nor univocal though, which makes it a future subject of study.
@@
Still with the goal of understanding the connection between the performance of a switching protocol with the properties of its sequencing scheme, we have introduced a structural plane composed of measurements of lacunarity and persistence analogous to the complexity-entropy plane used to classify data seriesRosso et al. (2007). Overall, we observe that the switching protocols are prone to become less performing as its persistence augments. This is understood by the fact that highly persistent strategies tend to look very much like an isolated game. Yet, there is the heterogeneity assessed by (log-)lacunarity, which can dramatically change the performance. For instance, the TM switching protocol outperforms both the RS and the periodical P[3,2]. These three cases have close log-lacunarity values, but significantly larger persistence one with respect to the other; however, when we compare the performance of P[4,1] with P[1,4], which have the same persistence, we understand that the increase of lacunarity is accompanied by a significant capital hike. A close picture is learned when we go from the Fb game to the periodic P[2,1] and therefrom to the TM game or farther afield to P[1,2]. As well-known, a strategy becomes Parrondian when the second game is played at the right time leading to a positive accumulation of capital; the lacunarity is a way to gauge that sparseness with respect to one of the games (i.e., game $A$ in our case), an insight that the cross-correlation is unable to provide us with. This explains the reason game P[1,4] has high lacunarity (wherein game $A$ is less frequent) than game P[4,1], for which game $A$ is rather frequent. In future work, we expect to explore the delicate balance between lacunarity and persistence from a analytical point of view as well as surveying the performance of these aperiodic series in history-dependent switching protocols.